OpenStudy (anonymous):

Angela ,Betty and Carol shared a sum of money in the ratio 4:2:5,if Carol received $15 more than Betty,how much money dis Angela receive?

OpenStudy (slaaibak):

suppose betty received x now carol received x + 15 so (x + 15)/x = 5/2 solving gives x = 10 so betty got 10, carol got 25 the ratio must be 4:2:5 which means angela got 20

OpenStudy (anonymous):

thanks a lot,we got stuck with this problem!